On the first day of the fun fair, there were 400 more boys than girls. On the second day, the number of boys increased by 30% and the number of girls increased by 70%. There were 2020 children on the second day. How many boys were there on the first day?
1
Expert's answer
2013-03-13T07:46:23-0400
On the first day of the fun fair, there were 400 more boys than girls. On
the second day, the number of boys increased by 30% and the number of girls
increased by 70%. There were 2020 children on the second day. How many boys
were there on the first day?
Let's denote the number of boys and girls on the first day by B and G respectivly. As there were 400 more boys than girls on the first day, then
B = G + 400.(1)
On the second day:
1.3B + 1.7G = 2020.(2)
We've got a system of equations. Let's solve it. Substituting (1) into (2) we obtain:
(G + 400)·1.3 + 1.7G = 2020 ==>
1.3G + 1.3*400 + 1.7G = 2020 ==>
3G = 1500 ==> G = 500 ==> B = 500 + 400 = 900.
Therefore, there were 500 girls and 900 boys on the first day.